# Hungarian Research Network (HUN-REN) Laboratory on Engineering and Management Intelligence Research Group of Operations Research and Decision Systems

### v12

@@ 4.2 The consequences of non-uniform distribution
In Group A, there are three home-away matches played by Russia and another team. Then the expected number of goals scored by the host Russia (denoted by $R$) equals
$\lambda_{Rj}^{(h)}=\left\{\begin{array}[]{ll}-5.42301\cdot W_{Rj}^{4}+15.49728\cdot W_{Rj}^{3}\\ -12.6499\cdot W_{Rj}^{2}+5.36198\cdot W_{Rj}+0.22862&\textrm{if }W_{Rj}\leq 0.9\\ \\ 231098.16153\cdot(W_{Rj}-0.9)^{4}-30953.10199\cdot(W_{Rj}-0.9)^{3}&\\ +1347.51495\cdot(W_{Rj}-0.9)^{2}-1.63074\cdot(W_{Rj}-0.9)+2.54747&\textrm{if }W_{Rj}>0.9,\end{array}\right.$
@@
The same simulation methodology has recently been applied in some studies on tournament design (Csató, 2022b, ; Csató,, 2023; Stronka,, 2024). Note that even if this Elo-based approach is not necessarily the best available simulation model, it is only used to compare the outcomes under the uniform and non-uniform draw. Thus, our assumptions seem to be adequate for comparative purposes as emphasised by Appleton, (1995).
@@
Figure 4 shows how the FIFA World Cup draw mechanism distorts the chances of winning the group and being the runner-up. The official rule (draw order 1-2-3-4) increases the probability of winning the group by more than 1 percentage point for the four UEFA teams in Pot 2 (9, 11, 12, 16) mostly at the expense of Mexico (14) and Serbia (25). Fortunately, these effects are somewhat mitigated by taking the likelihood of obtaining the second position into account. While draw order 1-2-4-3 does not differ much from the traditional order of 1-2-3-4, draw order 1-3-2-4 strongly reduces the distortions in the case of the above countries except for Serbia. Among the four draw orders, 4-2-3-1 is the less favourable for the weakest European team in Pot 2 (16) and the three UEFA members in Pot 3 (17, 18, 20).
1.51.5-1.5- 1.511-1- 10.50.5-0.5- 0.5000.50.50.50.511111.51.51.51.5RussiaGermanyBrazilPortugalArgentinaBelgiumPolandFranceSpainPeruSwitzerlandEnglandColombiaMexicoUruguayCroatiaDenmarkIcelandCosta RicaSwedenTunisiaEgyptSenegalIranSerbiaNigeriaAustraliaJapanMoroccoPanamaSouth KoreaSaudi ArabiaChange in percentage pointsAbsolute changes in the probability of qualificationOfficial (draw order 1-2-3-4)      Draw order 1-2-4-3Draw order 1-3-2-4                  Draw order 4-2-3-1
@@
Table 3: Summary statistics on the effects of four
draw procedures in the 2018 FIFA World Cup
Draw order 1-2-3-4 1-2-4-3 1-3-2-4 4-2-3-1
Change GW RU Q GW RU Q GW RU Q GW RU Q
$>1\%$ 4 0 1 3 0 0 0 0 1 1 0 1
$0.5\%$$1\%$ 2 2 2 3 2 2 3 0 2 1 2 5
$0.25\%$$0.5\%$ 1 4 8 1 5 4 2 4 5 7 6 8
$-0.25\%$$0.25\%$ 12 20 11 13 18 20 20 24 14 17 17 8
$-0.5\%$$-0.25\%$ 6 0 5 7 1 4 6 4 7 1 4 2
$-1\%$$-0.5\%$ 6 6 4 4 6 1 0 0 2 4 3 4
$<-1\%$ 1 0 1 1 0 1 1 0 1 1 0 4
  • Column GW shows the number of teams for which the change in the probability of winning the group is within the given interval.
  • Column RU shows the number of teams for which the change in the probability of being the runner-up is within the given interval.
  • Column Q shows the number of teams for which the change in the probability of qualification is within the given interval.
The probabilities of being the group winner or the runner-up are aggregated in Figure 5. Draw order 1-2-4-3 is somewhat better than the official 1-2-3-4: despite the increased impact on Costa Rica, it is less unfair for Russia, Croatia, Denmark, Iceland, and Sweden. Draw order 1-3-2-4 can be chosen if the bias for Mexico should be reduced, however, that is achieved at the expense of Serbia. Draw order 4-2-3-1 is not worth implementing because of the high distortions for several countries. These observations are reinforced by the summary statistics provided in Table 3.
44-4- 433-3- 322-2- 211-1- 1001111222233334444RussiaGermanyBrazilPortugalArgentinaBelgiumPolandFranceSpainPeruSwitzerlandEnglandColombiaMexicoUruguayCroatiaDenmarkIcelandCosta RicaSwedenTunisiaEgyptSenegalIranSerbiaNigeriaAustraliaJapanMoroccoPanamaSouth KoreaSaudi ArabiaChange in %Relative changes in the probability of qualificationOfficial (draw order 1-2-3-4)      Draw order 1-2-4-3Draw order 1-3-2-4                  Draw order 4-2-3-1
@@
Table 4: Summary statistics on the relative effects of
four draw procedures in the 2018 FIFA World Cup
@@
$>3\%$ 0 0 0 2
$2\%$$3\%$ 3 1 2 2
$1\%$$2\%$ 7 4 3 3
$-1\%$$1\%$ 16 24 21 17
$-2\%$$-1\%$ 4 2 4 4
$-3\%$$-2\%$ 2 1 1 3
$<-3\%$ 0 0 1 1
  • The table reports the number of national teams for which the change in the probability of qualification by the given draw order is within the appropriate interval.
Finally, Figure 6 uncovers the relative changes in the probability of qualification for the knockout stage as the same absolute distortions can be more costly for the weaker teams. In this respect, draw order 1-2-4-3 can be a reasonable alternative to the official 1-2-3-4 as it is less biased for almost all teams except for Costa Rica and Nigeria as shown by Table 4. The other two orders are especially unfavourable for Serbia and the three UEFA (draw order 1-2-4-3) or the three CAF (draw order 4-2-3-1) teams in Pot 3.
@@ 5 Discussion
Our paper has analysed the unfairness of the 2018 FIFA World Cup draw. First, the connection of the standard constrained group draw procedure—used by the FIFA and UEFA—to permutation generation has been presented. Second, we have examined how this procedure departs from a random draw among all feasible allocations and considered all alternatives by relabelling the pots. The official draw order (Pot 1, Pot 2, Pot 3, Pot 4) has turned out to be an optimal choice among the 24 alternative rules with respect to natural measures of fairness.
@@ Acknowledgements
The research was supported by the National Research, Development and Innovation Office under Grant FK 145838.

### v13

@@ 4.2 The consequences of non-uniform distribution
In Group A, three home-away matches are played by Russia and another team. Then the expected number of goals scored by the host Russia (denoted by $R$) equals
$\lambda_{Rj}^{(h)}=\left\{\begin{array}[]{ll}-5.42301\cdot W_{Rj}^{4}+15.49728\cdot W_{Rj}^{3}\\ -12.6499\cdot W_{Rj}^{2}+5.36198\cdot W_{Rj}+0.22863&\textrm{if }W_{Rj}\leq 0.9\\ \\ 231098.16153\cdot(W_{Rj}-0.9)^{4}-30953.10199\cdot(W_{Rj}-0.9)^{3}&\\ +1347.51495\cdot(W_{Rj}-0.9)^{2}-1.63074\cdot(W_{Rj}-0.9)+2.54747&\textrm{if }W_{Rj}>0.9,\end{array}\right.$
@@
The same simulation methodology has recently been applied in some studies on tournament design (Csató, 2022b, ; Csató,, 2023; Stronka,, 2024). Note that even if this Elo-based approach is not necessarily the best available simulation model, it is only used to compare the outcomes under a uniform and a non-uniform draw. Thus, our assumptions seem to be adequate for comparative purposes as emphasised by Appleton, (1995, p 534) with the following words: “Since our intention is to compare tournament designs, and not to estimate the chance of the best player winning any particular tournament, we may within reason take whatever model of determining winners that we please”.
Last but not least, a ranking should be established based on the results of group matches. According to the rules of the 2018 FIFA World Cup, the first tie-breaking criteria are as follows (FIFA,, 2018, Article 20.6): (1) higher number of points obtained in all group matches; (2) superior goal difference in all group matches; (3) higher number of goals scored in all group matches. In contrast to the official regulation, if some teams are still tied after considering (1)–(3), their ranking is decided by drawing of lots in our simulation.
Thus, a simulation run consists of the following steps for any draw mechanism (uniform or FIFA World Cup draw procedure with a given draw order):
  1. 1.
    The eight groups are drawn according to the draw mechanism used (1 million times);
  2. 2.
    The outcomes of all group matches are generated (10 times for each simulated draw);
  3. 3.
    The group rankings are determined.
The results for any team are the number of times it finishes on the first and the second place in its group, which are at most 10 million, respectively.
@@
Figure 4 shows how the FIFA World Cup draw mechanism distorts the chances of winning the group and being the runner-up. The official rule (draw order 1-2-3-4) increases the probability of winning the group by more than 1 percentage point for the four UEFA teams in Pot 2 (Spain [9], Switzerland [11], England [12], Croatia [16]) mostly at the expense of Mexico (14) and Serbia (25). Fortunately, these effects are somewhat mitigated by taking the likelihood of obtaining the second position into account. While the draw order 1-2-4-3 does not differ much from the traditional order of 1-2-3-4, the draw order 1-3-2-4 strongly reduces the distortions in the case of the above countries except for Serbia. Among the four draw orders, 4-2-3-1 is the less favourable for the weakest European team in Pot 2 (Croatia [16]) and the three UEFA members in Pot 3 (Denmark [17], Iceland [18], Sweden [20]).
1.51.5-1.5- 1.511-1- 10.50.5-0.5- 0.5000.50.50.50.511111.51.51.51.5Russia (1)Germany (2)Brazil (3)Portugal (4)Argentina (5)Belgium (6)Poland (7)France (8)Spain (9)Peru (10)Switzerland (11)England (12)Colombia (13)Mexico (14)Uruguay (15)Croatia (16)Denmark (17)Iceland (18)Costa Rica (19)Sweden (20)Tunisia (21)Egypt (22)Senegal (23)Iran (24)Serbia (25)Nigeria (26)Australia (27)Japan (28)Morocco (29)Panama (30)South Korea (31)Saudi Arabia (32)Change in percentage pointsAbsolute changes in the probability of qualificationOfficial (draw order 1-2-3-4)      Draw order 1-2-4-3Draw order 1-3-2-4                  Draw order 4-2-3-1
@@
Table 4: Summary statistics on the absolute changes of qualifying
probabilities by four draw procedures in the 2018 FIFA World Cup
Draw order 1-2-3-4 1-2-4-3 1-3-2-4 4-2-3-1 Change (pp) GW RU Q GW RU Q GW RU Q GW RU Q $>1$ 4 0 1 4 0 0 1 0 1 1 0 1 0.50.50.50.51111 2 2 2 2 2 2 2 0 2 1 3 5 0.250.250.250.250.50.50.50.5 1 7 6 1 5 4 2 3 4 7 4 8 0.250.25-0.25- 0.250.250.250.250.25 12 17 14 13 19 19 24 26 15 17 18 8 0.50.5-0.5- 0.50.250.25-0.25- 0.25 7 0 5 7 0 6 2 3 9 1 4 2 11-1- 10.50.5-0.5- 0.5 5 5 3 4 5 0 0 0 0 4 3 6 $<-1$ 1 1 1 1 1 1 1 0 1 1 0 2 Column GW/RU/Q shows the number of teams for which the absolute change in percentage points of the probability of winning the group/being the runner-up/qualification by the given draw order (the probability according to the official draw mechanism with the given draw order minus the probability according to a uniform draw) is within the given interval.
The probabilities of being the group winner or the runner-up are aggregated in Figure 5. The draw order 1-2-4-3 is somewhat better than the official 1-2-3-4: despite the increased impact on Costa Rica, it is less unfair for Russia, Croatia, Denmark, Iceland, and Sweden. The draw order 1-3-2-4 can be chosen if the bias for Mexico should be reduced, however, that is achieved at the expense of Serbia. The draw order 4-2-3-1 is not worth implementing because of the high distortions for several countries. These observations are reinforced by the summary statistics provided in Table 4.
44-4- 433-3- 322-2- 211-1- 100111122223333Russia (1)Germany (2)Brazil (3)Portugal (4)Argentina (5)Belgium (6)Poland (7)France (8)Spain (9)Peru (10)Switzerland (11)England (12)Colombia (13)Mexico (14)Uruguay (15)Croatia (16)Denmark (17)Iceland (18)Costa Rica (19)Sweden (20)Tunisia (21)Egypt (22)Senegal (23)Iran (24)Serbia (25)Nigeria (26)Australia (27)Japan (28)Morocco (29)Panama (30)South Korea (31)Saudi Arabia (32)Change in %Relative changes in the probability of qualificationOfficial (draw order 1-2-3-4)      Draw order 1-2-4-3Draw order 1-3-2-4                  Draw order 4-2-3-1
@@
Table 5: Summary statistics on the relative changes of qualifying
probabilities by four draw procedures in the 2018 FIFA World Cup
@@
$>3\%$ 0 0 0 1
$2\%$$3\%$ 3 1 3 4
$1\%$$2\%$ 6 4 2 4
$-1\%$$1\%$ 16 25 21 16
$-2\%$$-1\%$ 5 1 3 3
$-3\%$$-2\%$ 2 1 2 2
$<-3\%$ 0 0 1 2
  • The cells show the number of national teams for which the relative change in percentage of the probability of qualification by the given draw order (one minus (the probability according to the official draw mechanism with the given draw order divided by the probability according to a uniform draw)) is within the appropriate interval.
Finally, Figure 6 uncovers the relative changes in the probability of qualification for the knockout stage as the same absolute distortions can be more costly for the weaker teams. In this respect, the draw order 1-2-4-3 can be a reasonable alternative to the official 1-2-3-4 as it is less biased for almost all teams except for Costa Rica as shown by Table 5. The other two draw orders are especially unfavourable for Serbia and the three UEFA (draw order 4-2-3-1) or the three CAF (draw order 1-3-2-4) teams in Pot 3.
@@ 5 Discussion
Our paper has analysed the unfairness of the 2018 FIFA World Cup draw. First, the connection of the standard constrained group draw procedure—used by the FIBA, FIFA, and UEFA—to permutation generation has been presented. Second, we have examined how this procedure departs from a random draw among all feasible allocations and considered all alternatives by relabelling the pots. The official draw order (Pot 1, Pot 2, Pot 3, Pot 4) has turned out to be an optimal choice among the 24 alternative rules with respect to natural measures of fairness.
Even though constrained group draw is currently used only in basketball and football, draw constraints offer the unique solution to maximise the number of intercontinental games in a group stage. This can be advantageous in several tournaments; for instance, the 2023 World Men’s Handball Championship contained eight groups, one with two of the three African teams, and another with two of the four South American teams (Devriesere et al.,, 2024). Any governing body will likely adopt the procedure of FIFA if they introduce draw constraints.
@@ Acknowledgements
The comments of the reviewers Martin Becker, Florian Felice, and Michael A. Lapré have greatly improved the presentation of the results.
@@
The research was supported by the National Research, Development and Innovation Office under Grant FK 145838, and by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.